OpenStudy (anonymous):

Six years ago, I was half as old as i will be four years from now. How old will i be five years from now?

OpenStudy (anonymous):

can you please show me the solution in equation form plz

OpenStudy (anonymous):

let the present age be x six years ago = x-6 4 years from now = x+4 Accoding to question : x-6=1/2(x+4) x-6=(x+4)/2 2(x-6)=x+4 2x-12=x+4 2x-x=16 x=16 5 years from now = 16+5 = 21

hero (hero):

kush got it

OpenStudy (anonymous):

thanks hero

hero (hero):

Although kush, you left out an important part: 16+4 = 20

hero (hero):

Shows that 10 is half of 20

OpenStudy (anonymous):

yes yes sorry for that

OpenStudy (anonymous):

thnxs sooooo much both of u <3

OpenStudy (anonymous):

@ hero

OpenStudy (anonymous):

where did he miss tht step in the solution cuz i am a bit confused

OpenStudy (anonymous):

You don't actually need it for the solution. It's just checking that the calculated value satisfies the conditions stated in the problem. Always good to check your work.

OpenStudy (anonymous):

thnx so much :)